OpenStudy (anonymous):

solve the equation and round the result to nearest hundredth. -15x+73=26 7y-27=14

OpenStudy (anonymous):

-15x + 73 = 26 -15x + 73 - 73 = 26 - 73 combine like terms by getting the constants on the same side -15x = -47 divide by -15 x = 3.13

OpenStudy (anonymous):

Now you do y